OpenStudy (anonymous):

Find the probability (rolling 2 or 5) with one number cube. A. 1/2 B. 1/3 C.1/6 D.1

OpenStudy (anonymous):

1/3 or 2/6 two out of six possible numbers

OpenStudy (anonymous):

thx bahrom ...

OpenStudy (anonymous):

P ( 2) = 1/6 P ( 5) = 1/6 -> P(2) + P (5) = 2/6 = 1/3